OpenStudy (anonymous):

Use the quadratic formula to find the roots of the equation. 2x2 - 28x + 98 = 0 A. {-7, 7} B. {7} C. {-2, 49} D. no real number solution

hero (hero):

2(x^2 - 14x +49) = 0 2(x^2 - 7x - 7x + 49) = 0 2(x(x-7)-7(x-7)) = 0 2(x-7)(x-7) =0 2(x-7)^2 = 0 (x-7)^2 = 0 x-7 = 0 x = 7